Suppose a firm’s business operations are such that they mirror movements in the economy as a whole very closely; that is, the firm’s asset beta is 1. Find the equity beta for this firm for debt-equity ratios of 0, 1.9, 6.5, and 28. (Do not round intermediate calculations. Round your answers to 1 decimal place, e.g., 32.1.)
Answer #1

Asset beta = 1.0

If debt-equity ratio is 0.0:

Equity beta = Asset beta * [1 + Debt-equity ratio]
Equity beta = 1.0 * [1 + 0.0]
Equity beta = 1.00

If debt-equity ratio is 1.9:

Equity beta = Asset beta * [1 + Debt-equity ratio]
Equity beta = 1.00 * [1 + 1.9]
Equity beta = 1.00 * 2.9
Equity beta = 2.9

If debt-equity ratio is 6.5:

Equity beta = Asset beta * [1 + Debt-equity ratio]
Equity beta = 1.0 * [1 + 6.5]
Equity beta = 1.0 * 7.5
Equity beta = 7.5

If debt-equity ratio is 28.0:

Equity beta = Asset beta * [1 + Debt-equity ratio]
Equity beta = 1.0 * [1 + 28.0]
Equity beta = 1.0 * 29.0
Equity beta = 29.0